OpenStudy (anonymous):

Metal for help with a practice problem: What is the 4th term of the sequence? a1 = 8 and an = 6 + an – 1 a4 = _________

OpenStudy (solomonzelman):

\(\Large\color{black}{ a_n=6+a_{(n-1)} }\) so every time you add a 6 to previous term. Therefore, \(\Large\color{black}{ a_n=8+6(n-1) }\)

OpenStudy (solomonzelman):

(knowing that difference is +6, and the first term is 8)

OpenStudy (solomonzelman):

plug in 4 for n, to find the forth term. \(\Large\color{black}{ a_{\color{red}{4}}=8+6(\color{red}{4}-1) }\)

OpenStudy (solomonzelman):

you should be able to take it from here.